Will you please help me with this problem? For a holiday party, Laura is making a seven layer torte, which consists of seven 8 inch diameter layers with a cream topping on each layer. She knows that one recipe of cream will cover 55 square inches of cake. How many batches of the topping must she make to complete her torte? ( use pi=3.14 as an approx. value) please show work.
Answer by mananth(16946)

Area of topping = area of circle with diameter 8"
Radius = 4"
area = pi*r^2
pi^4^2
=pi*16
.
7 layers
= 7*pi*16
= 351.85 sq. in
.
one recipe covers 55 sq in
351.85 sqin will require 351.85/55
= 6.39
= 7 batches